# Deep Code Search with Naming-Agnostic Contrastive Multi-View Learning

### v1

@@ 4.2.1. Overall Performance
Table 1. MRR scores of each method on the three datasets.
Model CodeSearchNet CoSQA CoSQA-Var
OCoR 0.165 0.352 0.293
CSRS 0.231 0.373 0.231
DeepCS 0.164 0.472 0.433
CodeBERT 0.665 0.652 0.558
GraphCodeBERT 0.694 0.648 0.456
CoCLR 0.637 0.647 0.515
DGMS 0.579 0.471 0.399
NACS 0.701 0.708 0.704
Firstly, we investigate the overall code search performance of all methods on CodeSearNet, CoSQA and CoSQA-Var datasets. Table 1 shows the MRR scores of each code search model. From the results, we have the following observations:
@@
    Compared to OCoR, CSRS and DeepCS, pre-training based methods GraphCodeBERT, CoCLR and NACS achieve much better performance on all the three datasets, showing the effectiveness of the Bert-style pre-training method on improving code search performance.
@@
    Compared to all baselines, NACS achieves highest MRR scores and it consistently outperforms the best two baselines GraphCodeBert and CoCLR on the three datasets, showing that the robustness of NACS.
@@ 4.5. Impacts of Hyper-Parameters on Search Results (RQ4)
We investigate the impacts of $\lambda$ in Equation 17 on the performance of NACS. Figure 7 illustrates the performance of NACS on CoSQA and CoSQA-Var datasets when $\lambda$ is set to 0, 0.0001, 0.001, 0.01 and 0.1, respectively. We can observe that the best result are achieved when $\lambda$ is set to 0.0001 and larger $\lambda$ ($>0.1$) will significantly degrade the performance of NACS. Overall, it is easy to tune $\lambda$ as the best value falls in a relative narrow range. When $\lambda$ is set to 0, NACS is identical to CodeBERT. The performance of NACS declines from 0.708/0.704 to 0.652/0.558 when $\lambda$ is changed from 0.0001 to 0. The noticeable performance reduction shows that our proposed naming-agnostic contrastive multi-view code representation learning indeed improve the performance of code search.
@@
Figure 7. Impacts of $\lambda$ on the performance of NACS.
Figure 8. Impacts of $\tau$ on the performance of NACS.
5. Discussion and Future Work
@@
Even though NACS demonstrates promising performance on the code search task, there still exist a few potential threats that may affect the validity and the reliability of this work:
  • Scalability: Existing studies of code search all focus on the effectiveness of the model and pay little attention to the efficiency of code search and whether code search methods can scale well to a real large-scale codebase. This work also lacks the study of the scalability of code search, which affects the reliability of NACS in the real scenario of code search.
  • Robustness: The quantitative results are reported over the five runs of the models. However, the case studies in Section 4.2.2 only provide a few cases due to the page limit, which hinders readers from fully understanding the performance of NACS and all baselines in different cases, i.e., whether NACS is robust w.r.t. different queries.
  • Generalization to different programming languages: This study adopts three datasets that contain code snippets only written in Python, which may raise doubts about whether NACS can be generalized to other programming languages. As AST is the common method to model programming languages and our proposed method is applied to ASTs only, we believe NACS can be easily applied to other programming languages.
  • Evaluation metrics: This study adopts the most prevalent evaluation metric MRR on code search for evaluation. But there exists a threat that using one metric may not fairly evaluate the code search performance. Hence, we suggest considering more information retrieval evaluation metrics together to get a full picture of the performance of code search.
Considering the above limitations of this work, in the future, we plan to further explore the following directions to improve NACS:
  • More thorough evaluation of code search methods: We will collect more code search data containing real queries and corresponding code snippets written in more programming languages to construct large-scale evaluation benchmarks for evaluating the code search task. We will also construct subsets that contain diverse queries to better assess the robustness of code search methods when facing different queries. We will also adopt more information retrieval metrics in the study. Based on the new datasets and more evaluation metrics, the performance of NACS and other code search methods can be fully evaluated.
  • Explore lightweight architectures for code search: To reduce the cost of code search and improve the scalability of NACS, we will explore other designs of lightweight architectures to reduce both the training code and the search cost of NACS so that it can scale to a large-scale codebase.

### v2

@@ 4.2.1. Overall Performance
Table 1. MRR scores of each method on the four datasets.
Model CodeSearchNet-Python CodeSearchNet-Java CoSQA CoSQA-Var
OCoR 0.165 0.201 0.352 0.293
CSRS 0.231 0.386 0.373 0.231
DeepCS 0.164 0.199 0.472 0.433
CodeBERT 0.665 0.667 0.652 0.558
GraphCodeBERT 0.694 0.689 0.648 0.456
CoCLR 0.637 0.631 0.647 0.515
DGMS 0.579 0.555 0.471 0.399
NACS 0.701 0.705 0.708 0.704
Firstly, we investigate the overall code search performance of all methods on CodeSearchNet-Python, CodeSearchNet-Java, CoSQA and CoSQA-Var datasets. Table 1 shows the MRR scores of each code search model. From the results, we have the following observations:
@@
    Compared to OCoR, CSRS, DeepCS and DGMS, pre-training based methods CodeBERT, GraphCodeBERT, CoCLR and NACS achieve much better performance on all the four datasets, showing the effectiveness of the Bert-style pre-training method on improving code search performance.
@@
    Compared to all baselines, NACS achieves highest MRR scores and it consistently outperforms the best two baselines GraphCodeBert and CoCLR on the four datasets, showing that the robustness of NACS.
@@
  4. (4)
    NACS performs similarly on both CodeSearchNet-Python and CodeSearchNet-Java and it consistently outperforms baselines, showing the effectiveness of NACS on both Python and Java.
@@
Table 2. Running time of each method on CodeSearchNet-Java and CoSQA.
Model CodeSearchNet-Java CoSQA
Train Test Train Test
OCoR 284.968 hours 4.017 hours 742.260 min. 261.901 sec.
CSRS 14.638 hours 6.445 hours 37.254 min. 432.371 sec.
DeepCS 75.762 hours 10.932 hours 195.437 min. 737.075 sec.
CodeBERT 15.962 hours 2028.199 sec. 41.611 min. 8.815 sec.
GraphCodeBERT 33.977 hours 284.899 sec. 87.956 min. 10.715 sec.
CoCLR 55.364 hours 2583.092 sec. 109.104 min. 46.374 sec.
DGMS 45.191 hours 1553.489 sec. 115.999 min. 12.742 sec.
NACS 18.740 hours 531.692 sec. 48.529 min. 34.658 sec.
4.2.2. Running Time
We report the running time of each methods in Table 2. Since CodeSearchNet-Python and CodeSearchNet-Java have similar amounts of data and CoSQA and CoSQA-Var have the same amount of data, we only provide the training and the test time on CodeSearchNet-Java and CoSQA in Table 2. From the results, we have the following observations:
  1. (1)
    OCoR has the longest running time. The reason is that, during training, it counts the overlaps between identifiers in code and words in queries, which is a costly operation.
  2. (2)
    Pre-training based methods CodeBERT, GraphCodeBERT, CoCLR and NACS do not require much more training time compared to small models CSRS and DGMS. And their training time is even shorter in some cases. We can also observe that the test time of pre-training methods are generally much shorter compared to OCoR, CSRS, DeepCS and DGMS, showing that pre-training methods are efficient during inference and do not affect user experience on code search.
  3. (3)
    The running time of NACS is at the same level as the running time of other pre-training methods. On the largest dataset CodeSearchNet-Java (it is 25 times larger than CoSQA), its running time is acceptable compared to other baselines, showing the scalability of NACS on large codebases.
Overall, the running overhead of NACS is at the same level as other pre-training based methods and it shows strong scalability on large codebase.
@@ 4.2.3. Case Studies of Code Search Results
4.2.4. Relations between Learned Query Space and Code Space
Figure 7. Cosine similarity between query and corresponding ground-truth code snippet on the CoSQA dataset.
Code search models aim at matching a query and the ground-truth code snippet. The similarity between the query embedding and the code embedding can somehow demonstrate the ability of a code search model to match queries and corresponding ground-truth code snippets. We use the average embeddings ($\mathbf{t}$) of tokens in a query to represent the query and use the average embeddings ($\mathbf{c}$) of code tokens in the corresponding ground-truth code snippet to indicate the ground-truth code snippet. We calculate the average cosine similarity between $\mathbf{t}$ and $\mathbf{c}$ and visualize the results of GraphCodeBERT, CoCLR and NACS on CoSQA in Figure 7. From Figure 7, we can see that the average similarity between $\mathbf{t}$ and $\mathbf{c}$ for NACS is highest, showing that NACS is better at matching a query and the ground-truth code snippet.
@@ 4.5. Impacts of Hyper-Parameters on Search Results (RQ4)
Figure 8. Impacts of $\lambda$ on the performance of NACS.
Figure 9. Impacts of $\tau$ on the performance of NACS.
We investigate the impacts of $\lambda$ in Equation 17 on the performance of NACS. Figure 8 illustrates the performance of NACS on CoSQA and CoSQA-Var datasets when $\lambda$ is set to 0, 0.0001, 0.001, 0.01 and 0.1, respectively. We can observe that the best result are achieved when $\lambda$ is set to 0.0001 and larger $\lambda$ (>0.1>0.1) will significantly degrade the performance of NACS. Overall, it is easy to tune $\lambda$ as the best value falls in a relative narrow range. When $\lambda$ is set to 0, NACS is identical to CodeBERT. The performance of NACS declines from 0.708/0.704 to 0.652/0.558 when $\lambda$ is changed from 0.0001 to 0. The noticeable performance reduction shows that our proposed naming-agnostic contrastive multi-view code representation learning indeed improve the performance of code search.
@@
5. Conclusion
@@
In the future, we plan to explore other designs of lightweight architectures to reduce both the training code and the search cost of NACS to reduce the cost of code search and further improve the scalability of NACS. We also plan to collect more code search data containing real queries and corresponding code snippets written in more programming languages to construct large-scale evaluation benchmarks for better evaluating the code search task.
Acknowledgements.
This work was partially supported by Natural Science Foundation of Xiamen, China (No. 3502Z202471028), National Natural Science Foundation of China (No. 62002303, 42171456) and CCF-Tencent Open Fund (RAGR20210129).